OpenStudy (anonymous):

A rectangular barge 7 m long and 2 m wide floats in fresh water. How much will the barge sink when loaded with 600 kg of sand? Show all calculations leading to an answer.

OpenStudy (jtfan2000):

Water's density is ≈ 1,000kg/m³ The displaced water occupies about 600/1000=.6m³ The barge's area is 15m². V = L×W×H = .6m³ H (depth of sinking) = .6m³/15m² = .04m = 40mm”